9.) What volume of 0.750 M HCl is required to neutralize 10.0 mL of 1 M NaOH? Show your work. 10.) How many grams of CoHi20s are present in a 250 mL. solution with a concentration of 0.123 M? Show your work
Answer #1

9.

      HCl(aq) + NaOH(aq) -----------------> NaCl(aq) + H2O(l)

     1 mole       1 mole

HCl                                                                NaOH

M1 = 0.75M                                                 M2 = 1M

V1   =                                                            V2 = 10ml

n1 =1                                                            n2 = 1

                   M1V1/n1    =    M2V2/n2

                    V1              = M2V2n1/M1n2

                                      = 1*10*1/0.75*1   = 13.34ml

The volume of HCl is 13.34ml

10. molarity = 0.123M

    volume of solution = 250ml

molar mass of C6H12O6   = 180g/mole

molarity    = W*1000/G.M.Wt* volume of solution in ml

0.123        = W*1000/180*250

W             = 0.123*180*250/1000    = 5.535g

the mass of C6H12O6 is 5.535g